# Multi-group Uncertainty Quantification for Long-form Text Generation

### v1

@@ Abstract
While large language models are rapidly moving towards consumer-facing applications, they are often still prone to factual errors and hallucinations. In order to reduce the potential harms that may come from these errors, it is important for users to know to what extent they can trust an LLM when it makes a factual claim. To this end, we study the problem of uncertainty quantification of factual correctness in long-form natural language generation. Given some output from a large language model, we study both uncertainty at the level of individual claims contained within the output (via calibration) and uncertainty across the entire output itself (via conformal prediction). Moreover, we invoke multicalibration and multivalid conformal prediction to ensure that such uncertainty guarantees are valid both marginally and across distinct groups of prompts. Using the task of biography generation, we demonstrate empirically that having access to and making use of additional group attributes for each prompt improves both overall and group-wise performance. As the problems of calibration, conformal prediction, and their multi-group counterparts have not been extensively explored previously in the context of long-form text generation, we consider these empirical results to form a benchmark for this setting.
Multi-group Uncertainty Quantification for Long-form Text Generation
Terrance Liu Carnegie Mellon University Pittsburgh, PA 15213 terrancl@cs.cmu.edu                        Zhiwei Steven Wu Carnegie Mellon University Pittsburgh, PA 15213 zstevenwu@cmu.edu
@@ 1 Introduction
In recent years, researchers have developed stronger large language models that perform well on a variety of tasks across different domains (Touvron et al., 2023; Bubeck et al., 2023; Jiang et al., 2023; Anil et al., 2023). In particular, researchers have demonstrated the capabilities of such models by showcasing their ability to generate long-form text. However, as use of LLMs continues to grow, so do concerns over their tendency to hallucinate facts (Huang et al., 2023). As a result, there is a growing need for methods that can reduce hallucinations (Manakul et al., 2023; Zhang et al., 2023), perform abstention (Yang et al., 2023), or provide correctness guarantees (Kumar et al., 2023; Mohri and Hashimoto, 2024; Quach et al., 2023).
In our work, we focus on the latter—broadly speaking, uncertainty quantification of long-form large language model generations, which is a problem that has not previously been exhaustively explored. In contrast to the few existing works on factuality in long-form generations (Quach et al., 2023; Mohri and Hashimoto, 2024), we make the observation that while uncertainty guarantees may be valid under the full data distribution, they may not still be valid within individual subgroups of the distribution. For example, generations describing local politicians may be more prone to error than generations concerning national leaders. To address this unmet need, we introduce methods quantifying uncertainty in long-form text generation that are valid not only across a full distribution of prompts (i.e., marginally) but also across identifiable subgroups of prompts (i.e., conditionally).
@@
Concretely, given a set of claims produced by an LLM in response to some prompt, our goal is to provide a confidence score or uncertainty guarantee about the factual correctness of the output. We explore this problem in two settings. Given a set of claims contained within some long-form prompt response, we (1) ensure factuality at the individual claim level and (2) provide uncertainty guarantees across the whole set of claims. We approach problem (1) via (multi)calibration (Hébert-Johnson et al., 2018), in which one wishes to output a calibrated score for each claim, while for problem (2), we apply (multivalid) conformal prediction (Jung et al., 2022), selecting some subset of claims that with high probability are all factually correct.
@@
Empirically, we evaluate our methods using biography generation as a testbed for multi-group uncertainty quantification of factuality. As the problems of calibration, conformal prediction, and their multi-group counterparts have not been extensively explored previously in the context of long-form text generation, we consider these empirical results to form a benchmark for this setting. To induce subgroups, we collect an additional set of attributes about the public figure (i.e., entity) for whom the LLM is producing a biography. We note that these groups are intersecting, meaning that each entity can belong to multiple subgroups. Our empirical results demonstrate that for both problems (1) and (2), multicalibration and multivalid conformal prediction techniques improve measures of uncertainty relative to standard (marginal) calibration and conformal prediction methods. This advantage holds regardless of whether evaluation is conducted within groups or across the entire dataset.
@@ 1.1 Related work
The task of evaluating factuality for long-form generation is challenging: not only do generated outputs consist of many parts that must be scored individually, but also scoring each part requires prohibitively costly manual annotation. To help make evaluation more tractable, Min et al. (2023a) introduce FActSCORE, which converts any generation into a set of atomic facts/claims that are then labeled true or false. Using this evaluation metric, Min et al. (2023a) test LLMs’ ability to generate biographies and find that their generations are pervaded with errors.
@@
While the most natural method is to use a model’s output probabilities directly as a confidence (or uncertainty) score (Achiam et al., 2023), it has been shown that model probabilities are not-well calibrated (Guo et al., 2017). As a result, many works have recently proposed alternative methods for generating uncertainty scores that can then be used to refine or correct LLM outputs. Xiong et al. (2023), for example, explores whether a model can express uncertainty directly through prompting. Meanwhile, Wang et al. (2022) draw inspiration from self-consistency, in which a score is defined over how often an answer appears over multiple responses the same prompt. Our work is complementary to this line of work. Rather than proposing an entirely new uncertainty score function, we focus on how one can leverage existing scores to achieve uncertainty guarantees on LLM factuality.
@@
Similar to Detommaso et al. (2024), we study multicalibration for LLM outputs. However, while Detommaso et al. (2024) calibrate for correctness in question-answering, we are the first to apply multicalibration to claims decomposed from long-form text generation. In addition, our work also closely relates to that of Mohri and Hashimoto (2024), who aim to provide high probability guarantees of factuality in long-form generation. In particular, Mohri and Hashimoto (2024) frame this problem as a nested conformal prediction problem, applying methods from works like Min et al. (2023a) to produce sets of claims outputted by an LLM that achieve some marginally valid coverage. In other words, their method produces some generation that on average, contains a correct output with any user-specified probability. Our work, however, extends this problem to multivalid conformal prediction: we produce generations that are not only correct on average but are also conditionally correct across subgroups.
@@ 2 Preliminaries
In this section, we describe calibration and conformal prediction, as well as their multi-group guarantee versions, multicalibration and multivalid conformal prediction.
@@ 2.1 Calibration
Calibration, then, defines a simpler, more tractable condition, in which instead of ensuring guarantees across all possible values of $f$, we ensure a guarantee over coarser, level sets of our uncertainty score.
@@
Defining level sets is akin to dividing the output space of $f$ (i.e., $[0,1]$) into buckets. For example, one could round $f(X)$ to the nearest value in some predefined set of probabilities (e.g. $\{0,0.5,1.0\}$). One can view this definition of calibration as a desirable guarantee since it serves as a minimal condition for Equation 1—any $f$ that satisfies 1 must (at the very least) also be calibrated. We note that to evaluate calibration, we can consider the average squared calibration error (ASCE) of $f$.
@@ 2.1.1 Multicalibration
Multicalibration (Hébert-Johnson et al., 2018) was developed to provide accurate guarantees across overlapping subgroups (i.e., a sample can belong to many groups). Let $g:\mathcal{X}\to\{0,1\}$ be a group function that evaluates to $1$ if $X$ belongs to some group. We study, then, the setting in which there exists of set of groups $\mathcal{G}$ that corresponds to our data domain $\mathcal{D}$. We note that while the set of groups can be disjoint, the problem of multicalibration then becomes trivial in this case because one can simply split a dataset into disjoint sets that can then each be calibrated individually. Consequently, prior work typically considers the more interesting case where many intersecting groups comprise $\mathcal{G}$
Given a group function $g$, we define group average squared calibration error (gASCE) as the following:
@@ 2.2 Conformal Prediction
The second part of our work follows the problem statement outlined in Mohri and Hashimoto (2024). Unlike in calibration, where each claim contained in some long-form generation is treated individually, Mohri and Hashimoto (2024) instead define their problem in terms of pairs $(X,Y)$, where $X$ is some input prompt and $L(X)=Y\in\mathcal{Y}$ is the long-form generation outputted by a LLM $L$. Because $Y$ may or may not be supported by some reference ground truth $Y^{*}$,111e.g., in the case of FActScore (Min et al., 2023a), is $Y$ supported by Wikipedia. Mohri and Hashimoto (2024) define factuality in terms of entailment operations $Y^{*}\implies Y$. Furthermore, they rewrite this relation as $Y^{*}\in E(Y)=\{Y^{\prime}\in\mathcal{Y}:Y^{\prime}\implies Y\}$. This equivalent set notation, in other words, means that some reference ground truth $Y^{*}$ (e.g., Wikipedia article in Min et al. (2023a)) is contained in the set of possible texts $Y^{\prime}$ that support all claims made in the LLM output $Y$.
Given this notation, then, the goal is to find some uncertainty set $\mathcal{T}(L(X))$ s.t. $P_{\mathcal{D}}(Y\in\mathcal{T}(L(X)))=1-\alpha$. In the context of long-form text generation, this goal translates to taking as input the original LLM output $L(X)$ and producing a subset of claims $\mathcal{T}(L(X))$ such that with high probability, $1-\alpha$, all remaining claims are factually correct.
2.2.1 Multivalid Conformal Prediction
Similar to calibration, one may also desire group conditional coverage guarantees for intersecting groups. Known as multivalid conformal prediction (Jung et al., 2022), these guarantees are stronger than marginal conformal guarantees, holding also when conditioned on group membership. Using group functions $g$, as defined in Section 2.1.1, full multivalid coverage can be written as the following: Given some set of groups $\mathcal{G}$, we have that
@@ 2.2.1 Multivalid Conformal Prediction
for all group functions $g\in\mathcal{G}$.
@@ 3 Methods
Next, we introduce the methods (and their group-conditional variants) for applying calibration and conformal prediction to language model factuality. In particular, we organize these methods into two categories: (1) iterative "patching"-based algorithms and (2) linear regressor algorithms. As we mention previously, prior exploration of long-form text generation has been limited. While Mohri and Hashimoto (2024) evaluated one variant—split conformal (SC)—on a small set of entities, we are not aware of prior work that has considered other uncertainty quantification methods in this setting.
3.1 Iterative "patching" algorithms
The first category of algorithms can be characterized as "patching" algorithms. Given a base method for calibration or conformal prediction, one iterates through groups $g\in\mathcal{G}$ in which the method does poorly on. At each iteration, the patching algorithm corrects the bias (i.e., patches up the function) on just that subset of examples in the data domain (i.e., $g(x)=1$). Once some stopping condition is met, the final, "patched up" satisfies multi-group guarantees.
@@ 3.1 Iterative "patching" algorithms
For calibration, we consider the standard method of Histogram Binning (HB) (Zadrozny and Elkan, 2001), presented in Algorithm 1. This method, takes some base scoring function $f$ and discretizes the output space to a set of $p$-th level sets $S_{p}(f)$, as defined in Section 2.1. Specifically, given some target grid of values $p\in{[\frac{1}{m}]}$, we round $f$ to the closest value in this grid
@@
Algorithm 1 then applies a constant correction [^2] In Algorithms 1 and 2, we assume that you are given the true data distribution and therefore can calculate $\Delta_{p,g}$. In practice, as is done for our experiments, $\Delta_{p,g}$ is estimated using a calibration set. for each level set $S_{p}(f)$ in the grid, based on the calibration error of the model $f^{\prime}$.
In Algorithm 2, we present the multi-group version of histogram binning, known as Iterative Grouped Histogram Binning (IGHB) (Hébert-Johnson et al., 2018). In this algorithm, we instead apply a constant correction conditioned on $S_{p,g}$ (i.e., both the level set and group membership). At each step $t$, IGHB identifies $S_{p,g}$ for which the calibration error (weighted by the group size) is highest and then corrects it for this level set and group. The algorithm then continues until some stopping condition is met. In other words, the method iteratively patches the output $f^{\prime}$ for various groups $g\in\mathcal{G}$.
@@
We first present the Split Conformal (SC) method (Shafer and Vovk, 2008; Gupta et al., 2022). In particular, we consider the standard approach where one constructs a set of nested sets and each output set contains some subset $\mathcal{F(X)}_{t}$ of claims generated by the LLM.
Following Mohri and Hashimoto (2024), we define these nested sets $\mathcal{T}$ as thresholds sets where each set $\mathcal{F}(L(X))$ contains the set of all individual claims $\{x\in L(X)\mid f(x)>t\}$ for some scoring function $f$ [^3] The same scoring function $f$ that one may wish to (multi)calibrate can also be used for split conformal. More formally, we have that $\mathcal{F}(L(X))_{t\in\mathcal{T}}$ satisfies the nested sequence property if for $t,t^{\prime}\in\mathcal{T},t\leq t$, we have that $\mathcal{F}_{t}(L(X))\subseteq\mathcal{F}_{t^{\prime}}(L(X))$.
To construct these threshold sets, we have the score
@@
In Algorithm 3, we present the multivalid split conformal (MVSC) prediction technique that is closely related to methods originally proposed in Jung et al. (2022). Similar to IGHB, we start with some base threshold (i.e., the threshold $q_{\alpha}$ obtained from using split conformal). Then at each iteration $t$, we find the group $g_{t}$ that has the worst squared coverage error $\Delta_{t,g}$, weighted by the size of the group $P(g_{t}(X)=1)$. Then, we simply "patch" the thresholds for examples $\{(X,Y)\mid g_{t}(X)=1\}$, again using the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$th-quantile of scores for $(X,Y)$ beloing to group $g_{t}$. Like in IGHB, we continue patching the set of thresholds until some stopping criterion is met.

### v2

@@ Abstract
While past works have shown how uncertainty quantification can be applied to large language model (LLM) outputs, the question of whether resulting uncertainty guarantees still hold within sub-groupings of data remains open. In our work, given some long-form text generated by an LLM, we study uncertainty at both the level of individual claims contained within the output (via calibration) and across the entire output itself (via conformal prediction). Using biography generation as a testbed for this study, we derive a set of (demographic) attributes (e.g., whether some text describes a man or woman) for each generation to form such “subgroups” of data. We find that although canonical methods for both types of uncertainty quantification perform well when measuring across the entire dataset, such guarantees break down when examining particular subgroups. Having established this issue, we invoke group-conditional methods for uncertainty quantification—multicalibration and multivalid conformal prediction—and find that across a variety of approaches, additional subgroup information consistently improves calibration and conformal prediction within subgroups (while crucially retaining guarantees across the entire dataset). As the problems of calibration, conformal prediction, and their multi-group counterparts have not been extensively explored in the context of long-form text generation, we consider these results to form a benchmark for this setting.
@@ 1 Introduction
In recent years, researchers have developed stronger large language models that perform well on a variety of tasks across different domains [Touvron et al., 2023, Bubeck et al., 2023, Anil et al., 2023]. However, as use of LLMs continues to grow, so do concerns over their tendency to hallucinate facts [Huang et al., 2023]. As a result, there is a growing need for methods that can reduce hallucinations [Manakul et al., 2023, Zhang et al., 2023], perform abstention [Yang et al., 2023], or provide correctness guarantees [Kumar et al., 2023, Mohri and Hashimoto, 2024, Quach et al., 2023]. Our work focuses on the latter—broadly speaking, uncertainty quantification of long-form large language model generations.
@@
Concretely, given a set of claims produced by an LLM in response to some prompt, our goal is to provide a confidence score or uncertainty guarantee about the factual correctness of the output. We explore this problem in two settings: given a set of claims contained within some long-form prompt response, we (1) ensure factuality at the individual claim level and (2) provide uncertainty guarantees across the whole set of claims. We approach problem (1) via calibration, in which one wishes to output a calibrated score for each claim, while for problem (2), we apply conformal prediction [Shafer and Vovk, 2008], selecting a subset of claims that—with high probability—are all correct.
In contrast to existing works on uncertainty guarantees of long-form generations [Quach et al., 2023, Mohri and Hashimoto, 2024], we make the observation that while these guarantees may be valid under the full data distribution, they may not still be valid within individual subgroups of the distribution. For example, generations describing local politicians may be more prone to error than generations concerning national leaders. We choose biography generation as a testbed for multi-group uncertainty quantification, arguing that this problem is well-motivated, given that bias within biography generation has long been studied [De-Arteaga et al., 2019]. Having derived a set of subgroups using demographic information (e.g., whether an LLM output describes a man or woman), we find that when evaluated with respect to such groupings, canonical methods for calibration and conformal prediction indeed exhibit significant biases. [^1] Uncertainty can be epistemic and aleatoric, and sources of group bias can be categorized into either (or both) types. Our work, which focuses on atomic factuality in long-form generation, falls under epistemic uncertainty.
Having established such issues for standard uncertainty quantification approaches, we shift our attention to understanding to what extent such biases can be corrected. To address this unmet need, we introduce methods quantifying uncertainty in long-form text generation that are valid not only across a full distribution of prompts (i.e., marginally) but also across identifiable subgroups of prompts (i.e., conditionally). Invoking (1) multicalibration [Hébert-Johnson et al., 2018] and (2) multivalid conformal prediction [Jung et al., 2022], we categorize methods into two styles: iterative “patching” and linear regressor-based algorithms.
@@
Our results demonstrate that for both problems (1) and (2), multicalibration and multivalid conformal prediction techniques improve measures of uncertainty relative to standard (marginal) calibration and conformal prediction methods. This advantage holds regardless of whether evaluation is conducted within groups or across the entire dataset. As the problems of calibration, conformal prediction, and their multi-group counterparts have not been extensively explored in the context of long-form text generation, we consider these results to form a benchmark for this setting.
@@ 1.1 Related work
Evaluating factuality for long-form generation [Min et al., 2023a, Song et al., 2024, Wei et al., 2024, Bayat et al., 2024] is challenging: not only do generated outputs consist of many parts that must be scored individually, but also scoring each part requires prohibitively costly manual annotation. To make evaluation more tractable, Min et al. [2023a] introduce FActSCORE, which converts any generation into a set of atomic facts (claims) that are then labeled as true or false. Using this evaluation metric, Min et al. [2023a] test LLMs’ abilities to generate biographies and find that their generations are pervaded with errors.
@@
While a natural method for producing an uncertainty estimate is to use a model’s output probabilities directly as a confidence score [Achiam et al., 2023], it has been shown that model probabilities are not well calibrated [Guo et al., 2017]. As a result, many works have recently proposed alternative methods for generating uncertainty scores that can then be used to refine or correct LLM outputs [Wang et al., 2022, Xiong et al., 2023, Geng et al., 2024, Fadeeva et al., 2023, Vashurin et al., 2024]. We highlight that such work is complementary to our line of work—rather than proposing an entirely new uncertainty score function, we focus on how one can better leverage existing scores to produce uncertainty guarantees.
@@
We note that much of the prior work on multicalibration and multivalid conformal prediction are rooted in theory. Like Detommaso et al. [2024], our work tries to bridge the gap between theoretical insights and practical problems today (i.e., LLM generations). However, while Detommaso et al. [2024] calibrate for correctness in question-answering, we are the first to apply multicalibration to claims decomposed from long-form text generation. Moreover, unlike Detommaso et al. [2024], we consider uncertainty quantification in the form of conformal prediction.
In addition, our work closely relates to Mohri and Hashimoto [2024], which aims to provide high probability guarantees of factuality in long-form generation. In particular, Mohri and Hashimoto [2024] frame this problem as a nested conformal prediction problem, producing subsets of claims that achieve some marginally valid coverage (i.e., produce some generation that on average, contains a correct output with any user-specified probability). Our work, however, extends this problem to multivalid conformal prediction: we produce generations that are not only correct on average but are also conditionally correct across subgroups.
Finally, concurrent work by Cherian et al. [2024] also builds off this framework, but unlike our work, they introduce a new objective in which the goal is to instead guarantee that at least some given proportion of claims are retained. By applying a method proposed in Gibbs et al. [2023], Cherian et al. [2024]’s algorithm can (optionally) condition on group membership. However, their experiments include only 5 (non-overlapping) groups that are derived from the same feature, while our work focuses on the more challenging setting in which examples can simultaneously belong to many groups.
@@ 2.1 Calibration
Calibration, then, defines a simpler, more tractable condition, in which instead of ensuring guarantees across all possible values of $f$, it ensures a guarantee over coarser, level sets $S_{p}(f)$:
@@
Defining level sets is akin to dividing the output space of $f$ (i.e., $[0,1]$) into buckets. For example, one could round $f(X)$ to the nearest value in some predefined set of probabilities (e.g. $\{0,0.5,1.0\}$). One can view this definition of calibration as a desirable guarantee since it serves as a minimal condition for Equation 1—any $f$ that satisfies (1) must (at the very least) also be calibrated. We note that to evaluate calibration, we can consider the average squared calibration error (ASCE) of $f$.
@@
Multicalibration [Hébert-Johnson et al., 2018] was developed to provide accurate guarantees across overlapping subgroups (i.e., a sample can belong to many groups). Let $g:\mathcal{X}\to\{0,1\}$ be a group function that evaluates to $1$ if $X$ belongs to some group. We study, then, the setting in which there exists of set of groups $\mathcal{G}$ that corresponds to our data domain $\mathcal{D}$. While the set of groups can be disjoint, the problem of multicalibration then becomes trivial in this case because one can simply split a dataset into disjoint sets that can then each be calibrated individually. Consequently, prior work typically considers the more interesting case where many intersecting groups comprise $\mathcal{G}$.
Given a group function $g$, we define group average squared calibration error (gASCE) as:
@@ 2.2 Conformal Prediction
The second part of our work follows the problem statement outlined in Mohri and Hashimoto [2024]. Unlike in calibration, where each claim contained in some long-form generation is treated individually, Mohri and Hashimoto [2024] instead define their problem in terms of pairs $(X,Y)$, where $X$ is some input prompt and $L(X)=Y\in\mathcal{Y}$ is the long-form generation outputted by a LLM $L$. Because $Y$ may or may not be supported by some reference ground truth $Y^{*}$,222in the case of FActScore [Min et al., 2023a], ”is $Y$ supported by Wikipedia?” Mohri and Hashimoto [2024] define factuality in terms of entailment operations $Y^{*}\implies Y$. Furthermore, they rewrite this relation as $Y^{*}\in E(Y)=\{Y^{\prime}\in\mathcal{Y}:Y^{\prime}\implies Y\}$. This equivalent set notation, in other words, means that some reference ground truth $Y^{*}$ (e.g., a Wikipedia article in Min et al. [2023a]) is contained in the set of possible texts $Y^{\prime}$ that support all claims made in the LLM output $Y$.
Given this notation, the goal is to find some uncertainty set $\mathcal{T}(L(X))$ s.t. $P_{\mathcal{D}}(Y\in\mathcal{T}(L(X)))=1-\alpha$. In the context of long-form text generation, this goal translates to taking as input the original LLM output $L(X)$ and producing a subset of claims $\mathcal{T}(L(X))$ such that with high probability, $1-\alpha$, all remaining claims are factually correct.
We note that to empirically measure such guarantees, one can use the coverage error of $\mathcal{T}$ w.r.t the target error rate $\alpha$.
$|P_{\mathcal{D}}(Y\notin\mathcal{T}(X))-\alpha|$ (5)
Multivalid Conformal Prediction.
Similar to calibration, one may also desire group conditional coverage guarantees for intersecting groups. Known as multivalid conformal prediction [Jung et al., 2022], these guarantees are stronger than marginal conformal guarantees, holding also when conditioned on group membership. Using group functions $g$, as defined in Section 2.1, full multivalid coverage can be written as the following: Given some set of groups $\mathcal{G}$, we have that
@@
for all group functions $g\in\mathcal{G}$. Thus, target coverage guarantees $1-\alpha$ must hold both marginally and within all subgroups.
@@ 3 Methods
Next, we introduce the methods (and their group-conditional variants) for applying calibration and conformal prediction to language model factuality. We organize these methods into two categories: (1) iterative “patching”-based algorithms and (2) linear regressor algorithms. As mentioned previously, prior exploration of long-form text generation has been limited. While Mohri and Hashimoto [2024] evaluate one variant—split conformal (SC)—on a small set of entities, we are not aware of prior work that has considered other uncertainty quantification methods in this setting.
3.1 Iterative “patching” algorithms
The first category of algorithms can be characterized as patching algorithms. Given a base method for calibration or conformal prediction, one iterates through groups $g\in\mathcal{G}$ in which the method does poorly on. At each iteration, the algorithm corrects the bias (i.e., patches up the function) on just that subset of examples (i.e., $g(x)=1$). Once some stopping condition is met, [^3] In the standard formulation of iterative patching, the stopping criteria is set as a function over the number of bins so that one can prove guarantees about algorithm (see Roth [2022]). In practice, we found this stopping criteria to be too conservative, and so we instead run iterative patching on the calibration and test sets concurrently and use the calibration set to determine the stopping iteration (i.e., we enforce early stopping once we can no longer make improvements on the calibration set). the final, "patched up" function satisfies multi-group guarantees.
@@ 3.1 Iterative “patching” algorithms
For calibration, we consider Histogram Binning (HB) [Zadrozny and Elkan, 2001], presented in Algorithm 1. This method, takes some base scoring function $f$ and discretizes the output space to a set of $p$-th level sets $S_{p}(f)$, as defined in Section 2.1. Given some target grid of values $p\in{[\frac{1}{m}]}$, we round $f$ to the closest value in the grid
@@
Algorithm 1 then applies a constant correction [^4] In Algorithms 1 and 2, we assume true data distribution is given, and therefore we can calculate $\Delta_{p,g}$. In practice (and our experiments) $\Delta_{p,g}$ is estimated using a calibration set. for each level set $S_{p}(f)$ in the grid, based on the calibration error of the model $f^{\prime}$.
In Algorithm 2, we present the multi-group version of histogram binning, known as Iterative Grouped Histogram Binning (IGHB) [Hébert-Johnson et al., 2018]. In this algorithm, we instead apply a constant correction conditioned on $S_{p,g}$ (i.e., both the level set and group membership). At each step $t$, IGHB identifies $S_{p,g}$ for which the calibration error (weighted by the group size) is highest and then corrects it for this level set and group. The algorithm then continues until some stopping condition is met, iteratively patching $f^{\prime}$ for various groups $g\in\mathcal{G}$.
@@
We first present the Split Conformal (SC) method [Shafer and Vovk, 2008, Gupta et al., 2022]. In particular, we consider the standard approach where one constructs a set of nested sets and each output set contains some subset $\mathcal{F(X)}_{t}$ of claims generated by the LLM.
Following Mohri and Hashimoto [2024], we define these nested sets $\mathcal{T}$ as thresholds sets where each set $\mathcal{F}(L(X))$ contains the set of all individual claims $\{x\in L(X)\mid f(x)>t\}$ for some scoring function $f$. More formally, we have that $\mathcal{F}(L(X))_{t\in\mathcal{T}}$ satisfies the nested sequence property if for $t,t^{\prime}\in\mathcal{T},t\leq t^{\prime}$, we have that $\mathcal{F}_{t}(L(X))\subseteq\mathcal{F}_{t^{\prime}}(L(X))$.
To construct these threshold sets, we have that
@@
In Algorithm 3, we present the multivalid split conformal (MVSC) prediction technique that closely resembles methods originally proposed in Jung et al. [2022]. Similar to IGHB, we start with some base threshold (i.e., the threshold $q_{\alpha}$ obtained from using split conformal). Then at each iteration $t$, we find the group $g_{t}$ that has the worst squared coverage error $\Delta_{t,g}$, weighted by the size of the group $P(g_{t}(X)=1)$. Then, we simply "patch" the thresholds for examples $\{(X,Y)\mid g_{t}(X)=1\}$, again using the $\frac{\lceil(n+1)(1-\alpha)\rceil}{n}$th-quantile of scores for $(X,Y)$ belong to group $g_{t}$. Like in IGHB, we continue patching the set of thresholds until some stopping criterion is met.